# AttributionLab: Faithfulness of Feature Attribution Under Controllable Environments

### v1

@@ Abstract
Feature attribution explains neural network outputs by identifying relevant input features. How do we know if the identified features are indeed relevant to the network? This notion is referred to as faithfulness, an essential property that reflects the alignment between the identified (attributed) features and the features used by the model. One recent trend to test faithfulness is to design the data such that we know which input features are relevant to the label and then train a model on the designed data. Subsequently, the identified features are evaluated by comparing them with these designed ground truth features. However, this idea has the underlying assumption that the neural network learns to use all and only these designed features, while there is no guarantee that the learning process trains the network in this way. In this paper, we solve this missing link by explicitly designing the neural network by manually setting its weights, along with designing data, so we know precisely which input features in the dataset are relevant to the designed network. Thus, we can test faithfulness in AttributionLab, our designed synthetic environment, which serves as a sanity check and is effective in filtering out attribution methods. If an attribution method is not faithful in a simple controlled environment, it can be unreliable in more complex scenarios. Furthermore, the AttributionLab environment serves as a laboratory for controlled experiments through which we can study feature attribution methods, identify issues, and suggest potential improvements.
@@ 1 Introduction
Neural networks exhibit increasing capabilities as the scale of their design and their training data increases (Bubeck et al., 2023; Wei et al., 2022; Oquab et al., 2023; OpenAI, 2023; Caron et al., 2021; Brown et al., 2020). These capabilities are achieved through the use of basic architectural blocks  (Simonyan & Zisserman, 2015; He et al., 2016; Vaswani et al., 2017). Though we know the architectural design of these networks and know their computational graph explicitly, we do not have a human interpretable understanding of neural networks. One way to explain the neural network output is to identify important input features for a single prediction, an explanation paradigm known as input feature attribution. There has been an ongoing quest for finding attribution methods to explain neural network functions (Selvaraju et al., 2017; Zeiler & Fergus, 2014; Bach et al., 2015; Springenberg et al., 2015; Shrikumar et al., 2017; Fong et al., 2019; Lundberg & Lee, 2017; Zhang et al., 2021; Sundararajan et al., 2017; Schulz et al., 2020). However, one challenge remains: How can we know whether the features identified by attribution are aligned with features relevant to the neural network? I.e., how do we know if an attribution is faithful? An attribution may seem reasonable to us, but the neural network may use other input features. Conversely, an attribution may seem unreasonable but be faithful and indeed reflect the features relevant to the neural networks. Moreover, in the presence of multiple differing attribution explanations (Krishna et al., 2022; Khakzar et al., 2022), it is unclear which explanation to trust. The increasing complexity of networks, compounded by the rising complexity of feature attribution methodologies, further convolutes the problem.
@@
One recent trend to assess the faithfulness of attribution methods is through the use of synthetic data. Synthetic data are designed such that associations between features and labels are known to users (Arras et al., 2022; Agarwal et al., 2023; Zhou et al., 2022). However, as we discuss in Section 3 there is no guarantee that the learning process will train the network to use the designed associations in the dataset. Hence, the association learned by models can differ from the designed association in the synthetic dataset. Consequently, evaluation based on the designed association is not guaranteed to reflect the faithfulness of feature attribution methods. Furthermore, many evaluations usually report a performance score without any information on why an attribution method has limited performance. A test environment capable of uncovering properties and issues of methods can better contribute to the development of feature attribution research.
@@
In scientific experiments with complex setups and multiple variables, it is typical to use a laboratory setting to conduct controlled tests. Analogously, our work proposes a paradigm for providing a controlled laboratory environment. In this laboratory environment, both the neural networks and the datasets are designed such that we know which features are relevant to the network output. Thus, we obtain the ground truth attribution in this synthetic environment. We leverage this information for the faithfulness test by measuring the alignment between the ground truth attribution and attribution maps (Section 4). If an attribution method fails to pass the faithfulness test in the simulated setup, its performance in more complex scenarios can also be suboptimal. A controlled environment can also be used to study the behavior of attribution methods under various circumstances by adjusting or ablating variables to simulate different scenarios. With the help of proposed synthetic environments, we examine a broad range of attribution methods and investigate the impact of several crucial factors, including the choice of baseline and superpixel segmentation (Section 5). We make several observations from the test results and provide suggestions for improving their attribution performance. Furthermore, we show how the controlled environment can be used to analyze perturbation-based faithfulness evaluation metrics (Section 6).
2 Related work
Since the advent of deep neural networks (Krizhevsky et al., 2012; Simonyan & Zisserman, 2015; He et al., 2016), understanding how these complex models make predictions came under the spotlight (Simonyan et al., 2013). One approach is a simplified view of identifying features relevant to a prediction, known as feature attribution. Initial efforts to perform feature attribution focused on linear models (Simonyan et al., 2013; Bach et al., 2015) and backpropagation (Springenberg et al., 2015; Zeiler & Fergus, 2014). Subsequently, more principled approaches emerged, such as backpropagating output differences with respect to a reference input (Bach et al., 2015; Shrikumar et al., 2017) and axiomatic methods (Sundararajan et al., 2017; Lundberg & Lee, 2017) inspired by the Shapley value (Shapley, 1953). Meanwhile, intuitive approaches probing internal states (Selvaraju et al., 2017; Schulz et al., 2020; Chattopadhay et al., 2018) and optimizing input masks were introduced (Fong et al., 2019; Zhang et al., 2021). With these advancements, feature attribution evaluation became a central question. Literature has proposed sanity checks testing whether attribution changes upon randomizing the network’s weights (Adebayo et al., 2018) or if it is used on a different output class (Sixt et al., 2020). It also evaluates faithfulness by analyzing networks’s output when perturbing input features based on their relevance (Samek et al., 2016; Ancona et al., 2018; Montavon et al., 2018; Hooker et al., 2019; Rong et al., 2022), or show theoretically if they are aligned with axioms outlining desirable properties  (Sundararajan et al., 2017; Lundberg & Lee, 2017). Each perspective reveals different issues and properties (see Appendix B). A recent trend in evaluating faithfulness is designing datasets with known input-feature-output associations, enabling comparison between attribution and ground truth. Arras et al. (2022) generates datasets of shapes, Agarwal et al. (2023) generates synthetic graph datasets, and Zhou et al. (2022) proposes a dataset modification procedure to incorporate ground truth. However, these works do not guarantee that the network uses the intended ground truth associations. To address this problem, Khakzar et al. (2022) proposes a post-hoc solution of inverse feature generation by generating input features using the network. However, this direction faces challenges related to generated out-of-distribution input features. It is still unclear which parts of the generated features contribute and how they contribute.
3 A controllable environment to evaluate faithfulness
@@
To realize the controlled setup for evaluating explanations, one might design a dataset where the associations between input features and output labels are known: e.g., a simple dataset of squares and circles. Indeed, prior approaches (Arras et al., 2022; Zhou et al., 2022; Schuessler et al., 2021; Agarwal et al., 2023) design synthetic datasets and train a model on them. The underlying assumption is that a model learns the intended association in the dataset once the model achieves peak performance on the synthetic dataset during the learning process. However, it is unclear whether the trained network uses the entire square, straight lines, or just corners to identify the square. How can we know if a network uses an entire square for decision-making in this setup? This phenomenon is formulated as the Rashomom effect, which we rephrase in the following definition:
Figure 2: Designing data and model to set up a controllable environment for testing the faithfulness of attribution methods and analyzing their properties. To obtain the ground truth attribution, we explicitly design networks in tandem with inputs. The models follow conventional neural network designs and have sufficient complexity to evaluate various properties of attribution methods. The faithfulness test compares attribution results in the synthetic setting with the ground truth attribution.
@@
The Rashomon effect states that it is generally invalid to assume that a model with $100\%$ accuracy on a dataset $\mathcal{D}$ is ensured to learn the true labeling process. In reality, depending on the data distribution, we can have trained models that learn to ignore some features relevant to the labels. For instance, the model can learn other associations (e.g. use corners to classify a square). Specifically, neural networks tend to be over-parametrized for their learning tasks, hence making it harder to learn the true labeling function (Ba & Caruana, 2014; Frankle & Carbin, 2019). We provide empirical evidence for trained neural networks ignoring designed features (detailed setting in Appendix A). The result in Figure 1 shows a neural network, which achieves $100\%$ training accuracy but learns to solely use partial ground truth features designed for the dataset (objects at the edge). In a nutshell, a model can learn to perform correctly on the dataset but is not guaranteed to learn the intended ground truth features in the data. Therefore, we manually design and set the weights of the neural network.
By explicitly designing the data and the neural network according to specific properties, we know the ground truth features (features used by the network) in our synthetic setting. The knowledge of ground truth features can then be used for testing the faithfulness of feature attribution methods.
@@ 4 Design of data and neural network
We propose a modular setup where each component performs specific tasks and fulfills a purpose relevant to evaluating attribution methods. For the designs to facilitate the evaluation of feature attribution methods, we follow certain design principles. Firstly, the designs resemble real scenarios, such as image classification using a convolutional neural network (Figure 2). Designs that are similar to real-world cases can narrow the gap between real environments and synthetic environments. Hence, we can leverage synthetic environments to identify issues within attribution methods that are relevant to actual use cases. Furthermore, the design ensures that every ground-truth pixel is relevant and equally relevant. Specifically, with designed input data comprised of ground-truth (foreground) and baseline (background) pixels, designed neural networks have the following properties:
Proposition 1
@@
(Sensitivity property) The addition/removal of any ground-truth pixel to/from the background affects the output of the designed neural network.
Proposition 2
@@
(Symmetry property) The addition/removal of any ground-truth pixel to/from the background equally affects the output of the designed neural network.
@@
The Sensitivity property implies that in this designed model and designed dataset setup, every ground truth pixel is relevant to the neural network output. The symmetry property implies that every ground truth pixel in the synthetic dataset should have equal relevance to the output of the synthetic neural network. In fact, these properties are aligned with the sensitivity and symmetry feature attribution axioms  (Sundararajan & Najmi, 2020; Sundararajan et al., 2017; Lundberg & Lee, 2017), which axiomatically define the relevance of a feature.
@@
As our main environment, we design a setup that resembles real image classification tasks. The task is to identify the dominant color (in terms of number of pixels) within an input image (see Figure  2). The network first identifies predetermined colors and then counts the pixels for each. The designed dataset comprises color images, each containing $N_{C}$ patches of uniquely colored foreground objects and a distinct background color that is consistent across images, resulting in $N_{C}+1$ colors in every image. The images are input to the model in RGB format. By treating each foreground color as a class, we formulate a multi-class task to predict the dominant color in an image. The designed classification model outputs softmax scores using logits for $N_{C}$ classes, where the logit of each class corresponds to the sum of pixels of the respective color. The softmax operation on the neural network output makes pixels of colors other than the target negatively contribute to the output of the target color (after softmax). Thus the ground truth will have both positively and negatively contributing features, enabling us to evaluate whether an attribution method is able to discern between positive and negative contributions. The design of the model and the dataset also follow the properties in the following propositions (Proposition  1 and  2). Hence, we can infer that all pixels within the $N_{C}$ colored patches are relevant features.
@@ 4.1 Design details
Here, the number identification functions $I_{R}$, $I_{G}$, and $I_{B}$ each detect a predefined component of an RGB value and are implemented as shown in Equation 1. Hence, we have $C(r,g,b)=1$ for $r=R,g=G,b=B$, and $C(r,g,b)=0$ otherwise. The remaining $N_{R}$ redundant channels activate on any other colors not among the $N_{C}+1$ colors defined in our dataset. Specifically, if any pixel has a color that is neither a target color nor the background color, all $N_{R}$ redundant channels will activate at the position of this pixel. The activation mechanism of redundant channels is implemented as
@@
Consequently, $R(r,g,b)=1$ if all $C_{i}(r,g,b)=0$, and $R(r,g,b)=1$ if $C_{i}(r,g,b)=1$ for any $i$. Following the color detector, we have a CNN module that accumulates activation of the first $N_{C}$ channels respectively. We realize CNN accumulation in two settings, one with uniform weights in CNN, the other setting with non-uniform CNN weights. Figure 3 illustrates the working principle of pixel accumulation using CNN with non-uniform weights (more details in Appendix C.2). The remaining $N_{R}$ redundant channels have random connections to the input of this CNN module. Therefore, the CNN accumulation module will have unexpected outputs if the input images contain any color that is not seen in the training dataset. Lastly, we have an MLP that performs identity mapping, as the output of the CNN module already provides the logits of each target color for normal inputs. The rationale behind adding an identity MLP module is to preserve the conventional model architecture in image classification, where a model is typically designed as a CNN followed by an MLP head.
@@
We provide other synthetic settings in Appendix D. The incentive for having additional synthetic settings is to cover more use cases, for instance, gray input images, single output models, and other model weight design schemes (e.g., modulo computation).
5 Faithfulness test of attribution methods in AttributionLab

### v2

@@ Abstract
Feature attribution explains neural network outputs by identifying relevant input features. The attribution has to be faithful, meaning that the attributed features must mirror the input features that influence the output. One recent trend to test faithfulness is to fit a model on designed data with known relevant features and then compare attributions with ground truth input features. This idea assumes that the model learns to use all and only these designed features, for which there is no guarantee. In this paper, we solve this issue by designing the network and manually setting its weights, along with designing data. The setup, AttributionLab, serves as a sanity check for faithfulness: If an attribution method is not faithful in a controlled environment, it can be unreliable in the wild. The environment is also a laboratory for controlled experiments by which we can analyze attribution methods and suggest improvements.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Designing data and model to set up a controllable environment for testing the faithfulness of attribution methods and analyzing their properties. To obtain the ground truth attribution, we explicitly design networks in tandem with inputs. The models follow conventional neural network designs and have sufficient complexity (shown in Table 1). More synthetic environments, including different modules (e.g., a modulo computer) and different tasks (e.g., regression) are in the appendix C. The faithfulness test performs a sanity check on attribution results in the synthetic setting with the ground truth attribution.
Figure 2: Faithfulness test in AttributionLab. An attribution ($3^{rd}$ to the last image) is faithful if it is aligned with the Ground Truth attribution ($2^{nd}$ image) for a given input ($1^{st}$ image). When attributions are not faithful in this controlled environment, how can they be reliable in the wild? AttributionLab can serve as a sanity check for faithfulness and a tool to analyze current and future attribution methods. In the following, we also analyze different factors that cause the misalignment. The exact setting for this figure is in Appendix H.
@@
Neural networks exhibit increasing capabilities as the scale of their design and their training data increases (Caron et al., 2021; Wei et al., 2022; Oquab et al., 2023; OpenAI, 2023; Bubeck et al., 2023). These capabilities are achieved through the use of basic architectural blocks  (Krizhevsky et al., 2012; Vaswani et al., 2017). Though we know the architectural design of these networks and their computational graph explicitly, we do not have a human interpretable understanding of neural networks. One way to explain the neural network output is to identify important input features for a single prediction, an explanation paradigm known as input feature attribution. There has been an ongoing quest for finding attribution methods to explain neural network functions (Zeiler & Fergus, 2014; Selvaraju et al., 2017; Shrikumar et al., 2017; Sundararajan et al., 2017; Lundberg & Lee, 2017; Schulz et al., 2020; Zhang et al., 2021). However, one challenge remains: How can we know whether the attributed features are aligned with input features influencing the output of the neural network? I.e., how do we know if an attribution is faithful? An attribution may seem reasonable to us, but the neural network may use other input features (Ilyas et al., 2019). Conversely, an attribution may seem unreasonable but be faithful and indeed reflect the features relevant to the neural networks. In the presence of multiple differing attributions (Krishna et al., 2022; Khakzar et al., 2022), it is unclear which explanation to trust.
@@
One recent trend to assess faithfulness is using synthetic data. Synthetic data are designed such that associations between features and labels are known to users (Arras et al., 2022; Agarwal et al., 2023; Zhou et al., 2022). However, as we discuss in Section 2, there is no guarantee that the learning process will train the network to use the designed associations in the dataset. Hence, the association learned by models can differ from the designed association in the synthetic data. Thus, evaluation based on the designed association is not guaranteed to reflect faithfulness.
@@
In scientific experiments with complex setups and multiple variables, it is typical to use a laboratory setting to conduct controlled tests. Analogously, our work proposes a paradigm for providing a controlled laboratory environment. In this laboratory environment, both the neural networks and the datasets are designed such that we know which features are relevant to the network output. Thus, we obtain the ground truth attribution in this synthetic environment. We leverage this information for the faithfulness test by measuring the alignment between the ground truth attribution and attribution maps (Section 2 and 3). If an attribution method is not faithful in the simulated setup, its performance in more complex scenarios is unreliable. A controlled environment can also be used to study the behavior of attribution methods under various circumstances by adjusting or ablating variables to simulate different scenarios (Section 4).
Our contribution:
(1) We argue with empirical evidence and theoretical supports that designing only a synthetic dataset does not suffice to acquire ground truth attribution on a model, as a trained model can learn spurious features to achieve the task. (2) We propose a controllable environment consisting of a synthetic model and a synthetic dataset such that we obtain the ground truth attribution in this environment. A controllable environment can be used for our proposed faithfulness test as a form of sanity check for attribution methods and debugging attribution methods. We implement several controllable environments that cover various tasks. (3) We test mainstream attribution methods with our synthetic environment. Our experiment reveals several characteristics of these attribution methods. Based on the experiment results, we provide suggestions for future feature attribution research. (4) By simulating unseen data in AttributionLab we show that conventional perturbation analysis of faithfulness is unreliable.
2 Faithfulness test in controllable settings
@@
To realize the controlled setup for evaluating explanations, one might design a dataset where the associations between input features and output labels are known: e.g., a simple dataset of squares and circles. Indeed, prior approaches (Arras et al., 2022; Zhou et al., 2022; Schuessler et al., 2021; Agarwal et al., 2023) design synthetic datasets and train a model on them. The underlying assumption is that a model learns the intended association in the dataset once the model achieves peak performance on the synthetic dataset during the learning process. However, it is unclear whether the trained network uses the entire square, straight lines, or just corners to identify the square. How can we know if a network uses an entire square for decision-making in this setup? This phenomenon is formulated as the Rashomon effect, which we rephrase in the following definition:
@@
The Rashomon effect states that it is generally invalid to assume that a model with $100\%$ accuracy on a dataset $\mathcal{D}$ is ensured to learn the true labeling process. In reality, depending on the data distribution, we can have trained models that learn to ignore some features relevant to the labels. For instance, the model can learn other associations (e.g. use corners to classify a square). Specifically, neural networks tend to be over-parametrized for their learning tasks, hence making it harder to learn the true labeling function (Ba & Caruana, 2014; Frankle & Carbin, 2019). We provide empirical evidence for trained neural networks ignoring designed features (detailed setting in Appendix A). The result in Figure 3 shows a neural network, which achieves $100\%$ training accuracy but learns to solely use partial ground truth features designed for the dataset (objects at the edge). In a nutshell, a model can learn to perform correctly on the dataset but is not guaranteed to learn the intended ground truth features in the data. Furthermore, studies (Chen et al., 2019; Singh et al., 2020) demonstrate that attributions for a learned model are very fragile to spurious correlations (e.g. objects at the edge in above example). Therefore, we manually design and set the weights of the neural network.
By explicitly designing the data and the neural network according to specific properties, we know the ground truth features (features used by the network). Though designed neural networks can accurately provide ground truth attribution information, feature attribution evaluation only on designed neural networks might be biased. Hence, instead of using a synthetic environment for evaluation, we apply it for the faithfulness test as a form of sanity check for attribution methods. Attribution methods are usually designed to be model agnostic, that is, they do not assume a target model and can work in general. We concretize this claim with the following definition. Given a model class =$\mathcal{F}=\{f|f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}\}$, a dataset $D=\{d:d\in\mathbb{R}^{m+n}\}$, an attribution method $\mathcal{A}:\mathcal{F}\times\{D\}\rightarrow\mathbb{R}^{m}$, and a faithfulness measurement $M:\mathbb{R}^{m}\times\mathcal{F}\times\{D\}\rightarrow\mathbb{R}$, we define $(\gamma,\mathcal{F})$-model agnostic faithfulness as follows:
Definition 2.1.
$(\gamma,\mathcal{F})$-model agnostic faithfulness: An attribution method $\mathcal{A}$ is $(\gamma,\mathcal{F})$-model agnostic faithful, if
$\forall f\in\mathcal{F},M(\mathcal{A}(f,D),f,D)\geq\gamma.$ (1)
An attribution method that has $(\gamma,\mathcal{F})$-model agnostic faithfulness should have faithfulness measurement results over threshold $\gamma$ for all models belonging to model class $\mathcal{F}$. This is a mild definition, as it is defined over a set of limited models, while attribution methods also usually claim to work on a specific set of models. In addition, this definition still allows the attribution method’s faithfulness to vary in different models. However, $(\gamma,\mathcal{F})$-model agnostic faithfulness is hard to determine, as it would require validation through all possible models in the model class $\mathcal{F}$. Nevertheless, we can test if an attribution method is not $(\gamma,\mathcal{F})$-model agnostically faithful through proposed faithfulness test.
Faithfulness test:
An attribution method is not $(\gamma,\mathcal{F})$-model agnostic faithful, if
$\exists f\in\mathcal{F},M(\mathcal{A}(f,D),f,D)<\gamma.$ (2)
The faithfulness test only checks whether an attribution method is not model agnostic faithful, passing the test does not mean a method has model agnostic faithfulness. Nevertheless, this test provides an efficient and accurate means to identify underperforming attribution methods and ineffective settings for these methods. Previously, the difficulty lies in finding an accurate faithfulness measurement $M$. This is solved by our synthetic environment with designed $f$ and $D$. The main challenge remains in designing a model that can be considered to be in the model class $\mathcal{F}$. We discuss in the following section how we design synthetic models with conventional architecture and scale to satisfy this constraint.
@@ 3 Design of data and neural network
We propose a modular setup where each component performs specific tasks. For the designs to facilitate the faithfulness test on $(\epsilon,\mathcal{F})$-model agnostic faithfulness, we follow certain design principles. Firstly, the designs resemble real scenarios such as image classification using a convolutional neural network (Figure 1), to ensure that designed models belong to the desired model class. Furthermore, the design ensures that every ground-truth pixel is relevant and equally relevant. Specifically, with designed input data comprised of ground-truth (foreground) and baseline (background) pixels, designed neural networks have the following properties:
Proposition 3.1.
@@
(Sensitivity property) The addition/removal of any ground-truth pixel to/from the background affects the output of the model.
Proposition 3.2.
@@
(Symmetry property) The addition/removal of any ground-truth pixel to/from the background equally affects the output of the model.
@@
The Sensitivity property implies that in this designed model and dataset setup, every ground truth pixel is relevant to the neural network output. The symmetry property implies that every ground truth pixel in the synthetic dataset should have equal relevance to the output of the synthetic neural network. In fact, these properties are aligned with the sensitivity and symmetry feature attribution axioms  (Sundararajan & Najmi, 2020; Sundararajan et al., 2017; Lundberg & Lee, 2017), which axiomatically define the relevance of a feature.
@@
As our main environment, we design a setup that resembles real image classification tasks. The task is to identify the dominant color (in terms of number of pixels) within an input image (see Figure  1). The network first identifies predetermined colors and then counts the pixels for each. The designed dataset comprises color images, each containing $N_{C}$ patches of uniquely colored foreground objects and a distinct background color that is consistent across images, resulting in $N_{C}+1$ colors in every image. The images are input to the model in RGB format. By treating each foreground color as a class, we formulate a multi-class task to predict the dominant color in an image. The designed classification model outputs softmax scores using logits for $N_{C}$ classes, where the logit of each class corresponds to the sum of pixels of the respective color. The softmax operation on the neural network output makes pixels of colors other than the target negatively contribute to the output of the target color (after softmax). Thus the ground truth will have both positively and negatively contributing features, enabling us to test whether an attribution method is able to discern between positive and negative contributions. The design of the model and the dataset also follow the properties in the following propositions (Proposition  3.1 and  3.2). Hence, we can infer that all pixels within the $N_{C}$ colored patches are relevant features.
@@ 3.1 Model design details
Here, the number identification functions $I_{R}$, $I_{G}$, and $I_{B}$ each detect a predefined component of an RGB value and are implemented as shown in Equation 3. Hence, we have $C(r,g,b)=1$ for $r=R,g=G,b=B$, and $C(r,g,b)=0$ otherwise. The remaining $N_{R}$ redundant channels activate on any other colors not among the $N_{C}+1$ colors. Specifically, if any pixel has a color that is neither a target color nor the background color, all $N_{R}$ redundant channels will activate at the position of this pixel. The activation mechanism of redundant channels is implemented as
@@
Consequently, $R(r,g,b)=1$ if all $C_{i}(r,g,b)=0$, and $R(r,g,b)=0$ if $C_{i}(r,g,b)=1$ for any $i$. Following the color detector, we have a CNN module that accumulates activation of the first $N_{C}$ channels respectively. We realize CNN accumulation in two settings, one with uniform weights in CNN, the other setting with non-uniform CNN weights. Figure 4 illustrates the working principle of pixel accumulation using CNN with non-uniform weights (more details in Appendix C.2). The remaining $N_{R}$ redundant channels have random connections to the input of this CNN module. Therefore, the CNN accumulation module will have unexpected outputs if the input images contain any color that is not seen in the training dataset. Lastly, we have an MLP that performs identity mapping, as the output of the CNN module already provides the logits of each target color for normal inputs. The rationale behind adding an identity MLP module is to preserve the conventional model architecture in image classification, where a model is typically designed as a CNN followed by an MLP head.
Table 1: Model summary in terms of the number of layers and parameters. Ours refers to one of our designed models (detailed in Appendix E). All designed models in AttributionLab have adequate parameters and layers to simulate conventional CNN architectures.
Model # layers # parameters
Ours 17 675,033
ResNet-8 (He et al., 2016) 10 1,227,594
@@
We provide other synthetic settings in Appendix D. The incentive for having additional synthetic settings is to cover more use cases, for instance, gray input images, regression models, and other model weight design schemes (e.g., modulo computation).
4 Faithfulness test of attribution methods